# Linear Bandits with Partially Observable Features

### v2

@@ 6.2.2 Scenario 2
In Scenario 2, we set $d=d_{z}=2K$, implying that no latent features remain and allowing us to evaluate our algorithms under the condition $d>K$. As discussed in Section 4.1, we apply SVD to reduce the dimensionality of the observed features before constructing the augmented feature set. Figure 3 shows that our algorithms perform well even without unobserved features, both in terms of the level and robustness of the cumulative regret. Moreover, by incorporating dimensionality reduction, our algorithms remain effective even when the feature matrix is not full rank. Lastly, similar to Figure 2, cumulative regret in Case 2 converges slightly more slowly than in Case 1.
Meanwhile, the baseline OLB algorithms continue to struggle in identifying the optimal arm throughout the horizon. For LinUCB and LinTS, this behavior may be attributed to the curse of dimensionality, where regret scales linearly with the feature dimension $d$—a well-known limitations of linear bandit algorithms (Oswal et al., 2020; Tran et al., 2024). Additionally, since the true rewards are bounded by 1, the resulting small reward gaps between arms may further hinder the identification of the optimal arm. In the case of DRLasso, the similar behavior arises from the fixed feature setting: the algorithm uses an “averaged” context vector across the action space at each round for parameter estimation. Consequently, under the fixed feature setting, this strategy becomes effectively equivalent to using a single fixed vector throughout the entire learning horizon, thus limiting the expressiveness of the estimation.
@@ 7 Conclusion
In this work, we addressed a problem with partially observable features within the linear bandit framework. We showed that conventional algorithms ignoring unobserved features may suffer linear regret due to information loss, and introduced RoLF, a novel algorithm that accounts for latent features using only observed data without requiring prior knowledge. Our algorithm has achieved a tighter regret bound than existing methods, and this improvement is supported by our numerical experiments.
For the future work, from the perspective that our feature augmentation strategy reformulates the problem as another linear bandit problem without model misspecification, extending this approach to other reward models, e.g., generalized linear models, would be an interesting direction. Additionally, although we assumed the latent reward component to be linear in the unobserved features, we can relax such assumption by viewing the latent reward component as an exogenous factor that interferes with the learning process. This perspective allows for modeling the latent reward using a general function class without structural assumptions.
@@ Appendix A Related Works
In bandit problems, the learning agent learns only from the outcomes of chosen actions, leaving unchosen alternatives unknown (Robbins, 1952). This constraint requires a balance between exploring new actions and exploiting actions learned to be good, known as the exploration-exploitation tradeoff. Efficiently managing this tradeoff is crucial for guiding the agent towards the optimal policy. To address this, algorithms based on optimism in the face of uncertainty (Lai & Robbins, 1985) employ the Upper Confidence Bound (UCB) strategy, which encourages the learner to select actions with the highest sum of estimated reward and uncertainty. This approach adaptively balances exploration and exploitation, and has been widely used and studied in the context of linear bandits (Abe & Long, 1999; Auer, 2002; Dani et al., 2008; Rusmevichientong & Tsitsiklis, 2010). Notable examples include LinUCB (Li et al., 2010; Chu et al., 2011) and OFUL (Abbasi-yadkori et al., 2011), known for their practicality and performance guarantees. However, existing approaches differ from ours in two key aspects: (i) they assume that the learning agent can observe the entire feature vector related to the reward, and (ii) their algorithms have regret that scales linearly with the dimension of the observed feature vector, i.e., $\widetilde{O}(d\sqrt{T})$.
@@
In contrast, we develop an algorithm that achieves a sublinear regret bound by employing the doubly robust (DR) technique, thereby avoiding the linear dependence on the dimension of the feature vectors. The DR estimation in the framework of linear contextual bandits is first introduced by Kim & Paik (2019) and Dimakopoulou et al. (2019), and subsequent studies improve the regret bound in this problem setting by a factor of $\sqrt{d}$ (Kim et al., 2021, 2023b). A recent application (Kim et al., 2023c) achieves a regret bound of order $O(\sqrt{dT\log T})$ under IID features over rounds. However, the extension to non-stochastic or non-IID features remains an open question. To address this issue, we develop a novel analysis that applies the DR estimation to non-stochastic features, achieving a regret bound sublinear with respect to the dimension of the augmented feature vectors. Furthermore, we extend DR estimation to handle sparse parameters, thereby further improving the regret bound to be sublinear with respect to the reduced dimension.
@@
Our problem is more general and challenging than misspecified linear bandits, where the assumed reward model fails to accurately reflect the true reward, such as when the true reward function is non-linear (Lattimore & Szepesvári, 2020), or a deviation term is added to the reward model (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022). While our work assumes that the misspecified (or inaccessible) portion of the reward is linearly related to certain unobserved features, misspecified linear bandit problems can be reformulated as a special case of our framework. While the regret bounds in Lattimore & Szepesvári (2020), Bogunovic et al. (2021) and He et al. (2022) incorporate the sum of misspecification errors that may accumulate over the decision horizon, our work establishes a regret bound that is sublinear in the decision horizon $T$, not affected by misspecification errors. Ghosh et al. (2017) proposed a hypothesis test to decide between using linear bandits or MAB, demonstrating an $O(K\sqrt{T}\log T)$ regret bound when the total misspecification error exceeds $\Omega(d\sqrt{T})$. In contrast, our algorithm achieves an $O(\sqrt{(d+d_{h})T\log T})$ regret bound without requiring hypothesis tests for misspecification or partial observability.
@@
Last but not least, our problem appears similar to other literature addressing bandit problems with partially observable features (Tennenholtz et al., 2021; Park & Faradonbeh, 2022, 2024; Kim et al., 2023a; Zeng et al., 2025). In particular,Park & Faradonbeh (2022, 2024); Kim et al. (2023a) assume that the true features follow a specific distribution, typically Gaussian. Zeng et al. (2025) further assume that the true features evolve according to a linear dynamical system with additive Gaussian noise. Park & Faradonbeh (2022, 2024) and Zeng et al. (2025) construct the observed features as emissions from the true features via a known linear mapping, also corrupted by additive Gaussian noise, whereas Kim et al. (2023a) first corrupt the true features with Gaussian noise and then generate the observed features by masking elements of the corrupted features following an unknown Bernoulli distribution. In addition, all of these approaches aim to recover the true features: Park & Faradonbeh (2022, 2024) introduce a known decoder mapping from the observed features to the corresponding latent features; Kim et al. (2023a) leverage a Bayesian oracle strategy for estimation; and Zeng et al. (2025) estimate the true features using a Kalman filter. In contrast, our setting imposes no structural assumptions on either the observed or latent features, making the problem more general and challenging than those addressed in the indicated works. Furthermore, our approach does not attempt to recover any information related to latent features. Instead, we compensate for the lack of reward information due to unobserved features, in the sense that we project the inaccessible portion of the reward onto the orthogonal complement of the row space spanned by the observed feature vectors.
@@
On the other hand, Tennenholtz et al. (2021) assume that partially observed features are available as an offline dataset, and leverage the dataset to recover up to $L$ dimensions of the $d$-dimensional true features, where $L\leq d$ is the dimension of the partially observed features. This setting is different from ours in which partial observability arises naturally and no offline access is available. In their framework, the correlation between the observed and unobserved features is used to model the relationship between the estimator based on the observed features and that based on the true features, which requires further estimation of the unknown correlation. In contrast, our method is agnostic to such correlation, which makes our approach more practical. Moreover, we exploit the observed features via feature augmentation and DR estimation, resulting in faster regret convergence compared to their UCB-based approach.
@@ Appendix B Detailed Experimental Setup
For both scenarios, the features—including the true features $\mathbf{z}_{a}$, observed features $\mathbf{x}_{a}$, and unobserved features $\mathbf{u}_{a}$—are constructed differently based on the relationship between the row space spanned by the observed features, $\mathrm{R}(\mathbf{X})$, and that spanned by the unobserved features, $\mathrm{R}(\mathbf{U})$. In Case 1, the general case, the true features $\mathbf{z}_{a}$ for each arm $a\in[K]$ are sampled from $\mathcal{N}(\mathbf{0},\mathbf{I}_{d_{z}})$, and the observed features $\mathbf{x}_{a}$ are obtained by truncating the first $d$ elements of $\mathbf{z}_{a}$, following the definition given in Eq. 1.
@@
In Cases 2 and 3, on the other hand, the features are generated in a way to explicitly reflect the inclusion relationship between $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{U})$. Particularly, in Case 2, where $\mathrm{R}(\mathbf{U})\subseteq\mathrm{R}(\mathbf{X})$, the observed features $\mathbf{x}_{a}$, are sampled from $\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ for each $a\in[K]$. Then, we generate a coefficient matrix $\mathbf{C}_{\mathbf{u}}\in\mathbb{R}^{d_{u}\times d}$, where each element is sampled from $\mathrm{Unif}(-1,1)$, and construct $\mathbf{u}_{a}$ by computing $\mathbf{u}_{a}=\mathbf{C}_{\mathbf{u}}\mathbf{x}_{a}$. This construction ensures that $\mathrm{R}(\mathbf{U})$ lies within $\mathrm{R}(\mathbf{X})$. In Case 3, where $\mathrm{R}(\mathbf{X})\subseteq\mathrm{R}(\mathbf{U})$, we reverse the process by sampling $\mathbf{u}_{a}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d_{u}})$, generating $\mathbf{C}_{\mathbf{x}}\in\mathbb{R}^{d\times d_{u}}$ from $\mathrm{Unif}(-1,1)$, and setting $\mathbf{x}_{a}=\mathbf{C}_{\mathbf{x}}\mathbf{u}_{a}$, thereby ensuring $\mathrm{R}(\mathbf{X})\subseteq\mathrm{R}(\mathbf{U})$. In both Case 2 and Case 3, the true features $\mathbf{z}_{a}$ are formed by concatenating $\mathbf{x}_{a}$ and $\mathbf{u}_{a}$.
@@
After construction of the features, the orthogonal basis vectors $\{\mathbf{b}_{1}^{\top},\dots,\mathbf{b}_{K-d}^{\top}\}$ are derived via singular value decomposition (SVD) on the observed feature matrix $\mathbf{X}$, ensuring orthogonality to $\mathrm{R}(\mathbf{X})$ These basis vectors are linearly concatenated to $\mathbf{X}$ to form the augmented feature matrix. The reward parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$ is sampled from the uniform distribution $\mathrm{Unif}(-1/2,1/2)$, and the rewards are generated via dot products following the definition Eq. 2. We set the coupling probability $p$, a hyperparameter used in the sampling distribution of $\widetilde{a}_{t}$, is set to $0.6$ (see Eq. 9). The confidence parameter $\delta$, which is also a hyperparameter, is set to $10^{-4}$, and the total decision horizon is $T=1200$.
@@
Throughout the experiments, we fix the number of arms as $K=30$ and the dimensionality of the true features $d_{z}=35$. Furthermore, to accommodate both partial and full observability, we set $d_{z}\geq d$. Specifically, in Scenario 1, $d$ is set to $\lfloor d_{z}/2\rfloor=17$, indicating that only about half of the full feature space is observable to the agent. In Scenario 2, on the other hand, we set $d=2K=60$ and $d_{z}=d$, implying that latent features are absent. The setup of the second scenario also allows us to evaluate our algorithm under the condition where $d>K$. For each of the three structural cases (i.e., the relationships between $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{U})$) considered under both scenarios, we conduct five independent trials using different random seeds. The results are presented in terms of the sample mean and one standard deviation of the cumulative regret.
@@ Appendix C Robust to Latent Feature Algorithm with Ridge Estimator
12:        Sample $\widetilde{a}_{t}$ according to  Eq. 9.
@@
19:        Update $\widehat{\boldsymbol{\mu}}_{t}^{R}$ following Eq. 12 with $\widetilde{y}_{a,t}$ and update $\check{\boldsymbol{\mu}}_{t}^{R}$ following Eq. 13.
@@
$\widehat{\boldsymbol{\mu}}^{R}_{t}:=\left(\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}+I_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{y}_{a,\tau}\right),$ (12)
@@
For each $t$, let $\mathcal{E}_{t}\subseteq[t]$ denote an exploration phase such that for $\tau\in\mathcal{E}_{t}$ the action $a_{\tau}$ is sampled uniformly over $[K]$. Assume that $\|\boldsymbol{\mu}_{\star}\|_{\infty}\leq 1$ and $\|\widetilde{\mathbf{x}}_{a}\|_{\infty}\leq 1$ for all $a\in[K]$. Then for all round $t$ such that $|\mathcal{E}_{t}|\geq 32(1-p)^{-2}K^{2}\log(2Kt^{2}/\delta)$, with probability at least $1-3\delta$,
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}^{R}_{t}-{\boldsymbol{\mu}}_{\star})|\leq\frac{2}{\sqrt{t}}\left(\frac{\sigma}{p}\sqrt{K\log\frac{t+1}{\delta}}+\sqrt{K}\right).$
With $|\mathcal{E}_{t}|=O(K^{2}\log Kt)$ number of exploration, the DR Ridge estimator achieves $O(\sqrt{K/t})$ convergence rate over all $K$ rewards. This is possible because the DR pseudo-rewards defined in Eq. 10 impute the missing rewards for all arms $a\in[K]$ using $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$, based on the samples collected during the exploration phase, $\mathcal{E}_{t}$. With this convergence guarantee, we establish a regret bound for RoLF-Ridge, which is the adaptation of Algorithm 1 using the Ridge estimator.
@@
Suppose $\|\boldsymbol{\mu}_{\star}\|_{\infty}\leq 1$ and $\|\mathbf{z}_{a}\|_{\infty}\leq 1$ for all $a\in[K]$. For $\delta\in(0,1)$, with probability at least $1-4\delta$, the cumulative regret of the proposed algorithm using DR Ridge estimator is bounded by
$\mathrm{Reg}(T)\leq\frac{32K^{2}}{(1-p)^{2}}\log\frac{2dT^{2}}{\delta}+4\sqrt{2T\log\frac{2}{\delta}}+\frac{4\sqrt{T}}{K-1}+4\delta+8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{t+1}{\delta}}+1\right).$
The first and second terms come from the distribution of $a_{t}$ which is a combination of the $1-t^{-1/2}$-greedy policy and resampling up to $\rho_{t}:=\log((t+1)^{2}/\delta)/\log(1/p)$ trials. The third term is determined by the size of the exploration set, $\mathcal{E}_{t}$, while the last term arises from the estimation error bounded by the DR estimator as described in Theorem 4. The hyperparameter $p\in(1/2,1)$ balances the size of the exploration set in the third term and the estimation error in the last term. Overall, the regret is $O(\sqrt{KT\log T})$, which shows a significant improvement compared to the regret lower bound in Theorem 1 for any linear bandit algorithms that do not account for unobserved features and unobserved rewards.

### v3

@@ 6.2.2 Scenario 2
In Scenario 2, we set $d=d_{z}=2K$, implying that no latent features remain and allowing us to evaluate our algorithms under the condition $d>K$. As discussed in Section˜4.1, we apply SVD to reduce the dimensionality of the observed features before constructing the augmented feature set. Figure˜3 shows that our algorithms perform well even without unobserved features, both in terms of the level and robustness of the cumulative regret. Moreover, by incorporating dimensionality reduction, our algorithms remain effective even when the feature matrix is not full rank. Lastly, similar to Figure˜2, cumulative regret in Case 2 converges slightly more slowly than in Case 1.
Meanwhile, the baseline OLB algorithms continue to struggle in identifying the optimal arm throughout the horizon. For LinUCB and LinTS, this behavior may be attributed to the curse of dimensionality, where regret scales linearly with the feature dimension dd—a well-known limitation of linear bandit algorithms (Oswal et al., 2020; Tran et al., 2024). Additionally, since the true rewards are bounded by 1, the resulting small reward gaps between arms may further hinder the identification of the optimal arm. In the case of DRLasso, similar behavior arises from the fixed feature setting: the algorithm uses an “averaged” context vector across the action space at each round for parameter estimation. Consequently, under the fixed-feature setting, this strategy becomes effectively equivalent to using a single fixed vector throughout the entire learning horizon, thereby limiting the expressiveness of the estimation.
@@ 7 Conclusion
In this work, we addressed the problem of partially observable features within the linear bandit framework. We showed that conventional algorithms that ignore unobserved features may suffer linear regret due to information loss, and introduced RoLF, a novel algorithm that accounts for latent features using only observed data without requiring prior knowledge. Our algorithm achieves a tighter regret bound than existing methods, and this improvement is supported by our numerical experiments.
For future work, from the perspective that our feature augmentation strategy reformulates the problem as another linear bandit problem without model misspecification, extending this approach to other reward models, e.g., generalized linear models, would be an interesting direction. Additionally, although we assumed the latent reward component to be linear in the unobserved features, we can relax this assumption by viewing the latent reward component as an exogenous factor that interferes with the learning process. This perspective allows for modeling the latent reward using a general function class without structural assumptions.
@@ Appendix A Related Works
In bandit problems, the learning agent learns only from the outcomes of chosen actions, leaving unchosen alternatives unknown (Robbins, 1952). This constraint requires a balance between exploring new actions and exploiting actions learned to be good, known as the exploration-exploitation tradeoff. Efficiently managing this tradeoff is crucial for guiding the agent towards the optimal policy. To address this, algorithms based on optimism in the face of uncertainty (Lai & Robbins, 1985) employ the Upper Confidence Bound (UCB) strategy, which encourages the learner to select actions with the highest sum of estimated reward and uncertainty. This approach adaptively balances exploration and exploitation, and has been widely and studied in the context of linear bandits (Abe & Long, 1999; Auer, 2002; Dani et al., 2008; Rusmevichientong & Tsitsiklis, 2010). Notable examples include LinUCB (Li et al., 2010; Chu et al., 2011) and OFUL (Abbasi-yadkori et al., 2011), which are known for their practicality and performance guarantees. However, existing approaches differ from ours in two key aspects: (i) they assume that the learning agent can observe the entire feature vector related to the reward, and (ii) their algorithms have regret that scales linearly with the dimension of the observed feature vector, i.e., $\widetilde{O}(d\sqrt{T})$.
@@
In contrast, we develop an algorithm that achieves a sublinear regret bound by employing the doubly robust (DR) technique, thereby avoiding the linear dependence on the dimension of the feature vectors. The DR estimation in the framework of linear contextual bandits was first introduced by Kim & Paik (2019) and Dimakopoulou et al. (2019), and subsequent studies improve the regret bound in this problem setting by a factor of $\sqrt{d}$ (Kim et al., 2021, 2023b). A recent application (Kim et al., 2023c) achieves a regret bound of order $O(\sqrt{dT\log T})$ under IID features over rounds. However, the extension to non-stochastic or non-IID features remains an open question. To address this issue, we develop a novel analysis that applies the DR estimation to non-stochastic features, achieving a regret bound sublinear with respect to the dimension of the augmented feature vectors. Furthermore, we extend DR estimation to handle sparse parameters, thereby further improving the regret bound to be sublinear in the reduced dimension.
@@
Our problem is more general and challenging than misspecified linear bandits, where the assumed reward model fails to accurately reflect the true reward, such as when the true reward function is non-linear (Lattimore & Szepesvári, 2020), or when a deviation term is added to the reward model (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022). While our work assumes that the misspecified (or inaccessible) portion of the reward is linearly related to certain unobserved features, misspecified linear bandit problems can be reformulated as a special case of our framework. While the regret bounds in Lattimore & Szepesvári (2020), Bogunovic et al. (2021) and He et al. (2022) incorporate the sum of misspecification errors that may accumulate over the decision horizon, our work establishes a regret bound that is sublinear in the decision horizon $T$ and is not affected by misspecification errors. Ghosh et al. (2017) proposed a hypothesis test to decide between using linear bandits or MAB, demonstrating an $O(K\sqrt{T}\log T)$ regret bound when the total misspecification error exceeds $\Omega(d\sqrt{T})$. In contrast, our algorithm achieves an $O(\sqrt{(d+d_{h})T\log T})$ regret bound without requiring hypothesis tests for misspecification or partial observability.
@@
Last but not least, our problem appears similar to the literature addressing bandit problems with partially observable features (Tennenholtz et al., 2021; Park & Faradonbeh, 2022, 2024; Kim et al., 2023a; Zeng et al., 2025). In particular, Park & Faradonbeh (2022, 2024); Kim et al. (2023a) assume that the true features follow a specific distribution, typically Gaussian. Zeng et al. (2025) further assume that the true features evolve according to a linear dynamical system with additive Gaussian noise. Park & Faradonbeh (2022, 2024) and Zeng et al. (2025) construct the observed features as emissions from the true features via a known linear mapping, also corrupted by additive Gaussian noise, whereas Kim et al. (2023a) first corrupt the true features with Gaussian noise and then generate the observed features by masking elements of the corrupted features following an unknown Bernoulli distribution. In addition, all of these approaches aim to recover the true features: Park & Faradonbeh (2022, 2024) introduce a known decoder mapping from the observed features to the corresponding latent features; Kim et al. (2023a) leverage a Bayesian oracle strategy for estimation; and Zeng et al. (2025) estimate the true features using a Kalman filter. In contrast, our setting imposes no structural assumptions on either the observed or latent features, making the problem more general and challenging than those addressed in the aforementioned works. Furthermore, our approach does not attempt to recover any information related to latent features. Instead, we compensate for the lack of reward information due to unobserved features, in the sense that we project the inaccessible portion of the reward onto the orthogonal complement of the row space spanned by the observed feature vectors.
@@
On the other hand, Tennenholtz et al. (2021) assume that partially observed features are available as an offline dataset, and leverage the dataset to recover up to $L$ dimensions of the $d$-dimensional true features, where $L\leq d$ is the dimension of the partially observed features. This setting is different from ours in which partial observability arises naturally and no offline access is available. In their framework, the correlation between the observed and unobserved features is used to model the relationship between the estimator based on the observed features and that based on the true features, which requires further estimation of the unknown correlation. In contrast, our method is agnostic to such correlation, which makes our approach more practical. Moreover, we exploit the observed features via feature augmentation and DR estimation, resulting in a faster convergence rate of regret compared to their UCB-based approach.
@@ Appendix B Detailed Experimental Setup
For both scenarios, the features—including the true features $\mathbf{z}_{a}$, observed features $\mathbf{x}_{a}$, and unobserved features $\mathbf{u}_{a}$—are constructed differently based on the relationship between the row space spanned by the observed features, $\mathrm{R}(\mathbf{X})$, and the row space spanned by the unobserved features, $\mathrm{R}(\mathbf{U})$. In Case 1, the general case, the true features $\mathbf{z}_{a}$ for each arm $a\in[K]$ are sampled from $\mathcal{N}(\mathbf{0},\mathbf{I}_{d_{z}})$, and the observed features $\mathbf{x}_{a}$ are obtained by truncating the first $d$ elements of $\mathbf{z}_{a}$, following the definition given in Eq.˜1.
@@
In Cases 2 and 3, on the other hand, the features are generated in a way to explicitly reflect the inclusion relationship between $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{U})$. Particularly, in Case 2, where $\mathrm{R}(\mathbf{U})\subseteq\mathrm{R}(\mathbf{X})$, the observed features $\mathbf{x}_{a}$ are sampled from $\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ for each $a\in[K]$. Then, we generate a coefficient matrix $\mathbf{C}_{\mathbf{u}}\in\mathbb{R}^{d_{u}\times d}$, where each element is sampled from $\mathrm{Unif}(-1,1)$, and construct $\mathbf{u}_{a}$ by computing $\mathbf{u}_{a}=\mathbf{C}_{\mathbf{u}}\mathbf{x}_{a}$. This construction ensures that $\mathrm{R}(\mathbf{U})$ lies within $\mathrm{R}(\mathbf{X})$. In Case 3, where $\mathrm{R}(\mathbf{X})\subseteq\mathrm{R}(\mathbf{U})$, we reverse the process by sampling $\mathbf{u}_{a}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d_{u}})$, generating $\mathbf{C}_{\mathbf{x}}\in\mathbb{R}^{d\times d_{u}}$ from $\mathrm{Unif}(-1,1)$, and we set $\mathbf{x}_{a}=\mathbf{C}_{\mathbf{x}}\mathbf{u}_{a}$, thereby ensuring $\mathrm{R}(\mathbf{X})\subseteq\mathrm{R}(\mathbf{U})$. In both Case 2 and Case 3, the true features $\mathbf{z}_{a}$ are formed by concatenating $\mathbf{x}_{a}$ and $\mathbf{u}_{a}$.
@@
After the construction of the features, the orthogonal basis vectors $\{\mathbf{b}_{1}^{\top},\dots,\mathbf{b}_{K-d}^{\top}\}$ are derived via singular value decomposition (SVD) on the observed feature matrix $\mathbf{X}$, ensuring orthogonality to $\mathrm{R}(\mathbf{X})$. These basis vectors are linearly concatenated to $\mathbf{X}$ to form the augmented feature matrix. The reward parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$ is sampled from the uniform distribution $\mathrm{Unif}(-1/2,1/2)$, and the rewards are generated via dot products following the definition Eq.˜2. The coupling probability $p$, a hyperparameter used in the sampling distribution of $\widetilde{a}_{t}$, is set to $0.6$ (see Eq.˜9). The confidence parameter $\delta$, which is also a hyperparameter, is set to $10^{-4}$, and the total decision horizon is $T=1200$.
@@
Throughout the experiments, we fix the number of arms at $K=30$ and the dimensionality of the true features $d_{z}=35$. Furthermore, to accommodate both partial and full observability, we set $d_{z}\geq d$. Specifically, in Scenario 1, $d$ is set to $\lfloor d_{z}/2\rfloor=17$, indicating that only about half of the full feature space is observable to the agent. In Scenario 2, on the other hand, we set $d=2K=60$ and $d_{z}=d$, implying that latent features are absent. The setup of the second scenario also allows us to evaluate our algorithm in the case where $d>K$. For each of the three structural cases (i.e., the relationships between $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{U})$) considered under both scenarios, we conduct five independent trials using different random seeds. The results are presented in terms of the sample mean and one standard deviation of the cumulative regret.
@@ Appendix C Robust to Latent Feature Algorithm with Ridge Estimator
12:   Sample $\widetilde{a}_{t}$ according to  Eq.˜9.
@@
19:   Update $\widehat{\boldsymbol{\mu}}_{t}^{R}$ following Eq.˜12 with $\widetilde{y}_{a,t}$ and update $\check{\boldsymbol{\mu}}_{t}^{R}$ following Eq.˜13.
@@
$\widehat{\boldsymbol{\mu}}^{R}_{t}:=\left(\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{y}_{a,\tau}\right),$ (12)
@@
For each $t$, let $\mathcal{E}_{t}\subseteq[t]$ denote an exploration phase such that for each $\tau\in\mathcal{E}_{t}$ the action $a_{\tau}$ is sampled uniformly over $[K]$. Assume that $\|\boldsymbol{\mu}_{\star}\|_{\infty}\leq 1$ and $\|\widetilde{\mathbf{x}}_{a}\|_{\infty}\leq 1$ for all $a\in[K]$. Then with probability at least $1-3\delta$,
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}^{R}_{t}-{\boldsymbol{\mu}}_{\star})|\leq\frac{2}{\sqrt{t}}\left(\frac{\sigma}{p}\sqrt{K\log\frac{t+1}{\delta}}+\sqrt{K}\right),$
for all rounds $t$ such that $|\mathcal{E}_{t}|\geq 32(1-p)^{-2}K^{2}\log(2Kt^{2}/\delta)$.
With $|\mathcal{E}_{t}|=O(K^{2}\log Kt)$ exploration rounds, the DR Ridge estimator achieves an $O(\sqrt{K/t})$ convergence rate over all $K$ rewards. This is possible because the DR pseudo-rewards defined in Eq.˜10 impute the missing rewards for all arms $a\in[K]$ using $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$, based on the samples collected during the exploration phase, $\mathcal{E}_{t}$. With this convergence guarantee, we establish a regret bound for RoLF-Ridge, which is an adaptation of Algorithm˜1 using the Ridge estimator.
@@
Suppose $\|\boldsymbol{\mu}_{\star}\|_{\infty}\leq 1$ and $\|\mathbf{z}_{a}\|_{\infty}\leq 1$ for all $a\in[K]$. For $\delta\in(0,1)$, with probability at least $1-4\delta$, the cumulative regret of the proposed algorithm using the DR Ridge estimator is bounded by
$\mathrm{Reg}(T)\leq\frac{32K^{2}}{(1-p)^{2}}\log\frac{2dT^{2}}{\delta}+2\sqrt{2T\log\frac{2}{\delta}}+\frac{4\sqrt{T}}{K-1}+4\delta+8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{T}{\delta}}+1\right),$
The first and second terms come from the distribution of $a_{t}$, which is a combination of the $1-t^{-1/2}$-greedy policy and resampling up to $\rho_{t}:=\log((t+1)^{2}/\delta)/\log(1/p)$ trials. The third term is determined by the size of the exploration set, $\mathcal{E}_{t}$, while the last term arises from the estimation error bounded by the DR estimator as described in Theorem˜4. The hyperparameter $p\in(1/2,1)$ balances the size of the exploration set in the third term and the estimation error in the last term. Overall, the regret is $O(\sqrt{KT\log T})$, which shows a significant improvement compared to the regret lower bound in Theorem˜1 for any linear bandit algorithm that does not account for unobserved features and unobserved rewards.